OpenStudy (anonymous):

solve for x: 3x^2 + 9x +4=0 please show all work...thanks

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

so then x= -0.5425 and x= -2.4575

OpenStudy (anonymous):

Apply the method of completing the squares

hartnn (hartnn):

your answer is correct @stardust3829

OpenStudy (anonymous):

3x^2 + 9x +4=0 Divide through by 3: x^2 + 3x + 4/3 = 0 x^2 + 3x = -4/3 x^2 + (3/2)x + (3/2)^2 = -4/3 + (3/2)^2 (x + 3/2)^2 = -4/3 + 9/4 (x + 3/2)^2 = 11/12 (x + 3/2) = sqrt(11/12) x = - 3/2 (+/-) sqrt(11/12)

OpenStudy (anonymous):

This will give x = -0.5426, -2.4574